OpenStudy (kpop4life123):

I WILL FAN AND MEDAL!!!!!!!!!!! If (5^4)m = 5^12, what is the value of m?

Nnesha (nnesha):

\[\huge\rm (x^m)^n=x^{m \times n}\] so \[\huge\rm (5^4)^m=5^{4 \times m}=5^{12}\]what would you multiply with 4 to get 12